OpenStudy (anonymous):

how do you get the answer 13/420 from 13/126-13/180?

OpenStudy (anonymous):

its a bit tricky but you can get that by giving them a common denominator then simplifying the operation. first to set them to a common denominator, multiply each by the others denominator 13/126 * 180/180 = 2340/22680 12/180 * 126/126 = 1512/22680 Note that both 180/180 and 126/126 reduce to 1 so you arent actually changing the math. Then just perform the subtraction as normal 2340/22680-1512/22680 = 702/22680 This looks really ugly but if you simplify it by dividing out common factors (you can divide each by 54) you end up with 13/420

OpenStudy (anonymous):

awesome! thank you so much :) @LifeEngineer